OpenStudy (anonymous):

Find the surface area for a sphere with a radius of 10 feet. @iGreen

Parth (parthkohli):

How difficult can it be once you know the formula?

OpenStudy (aaronandyson):

\[S.A = 4 \times \pi \times r^2\]

OpenStudy (aaronandyson):

so now

OpenStudy (aaronandyson):

\[4 \times \pi \times 10 ^2\]

OpenStudy (aaronandyson):

\[4 \times \pi \times 10 \times 10\]

OpenStudy (aaronandyson):

10^2 = 10 times 10

OpenStudy (aaronandyson):

\[4 \times \pi \times 100\]

OpenStudy (aaronandyson):

\[400 \times \pi\]

OpenStudy (aaronandyson):

pi can either be taken as \[\frac{ 22 }{ 7 } or 3.14\]

OpenStudy (anonymous):

Thank you!